Given the following inverse demand and supply equations:

P = 24 - 6 Q

P = 4 Q

1. What is the before tax equilibrium quantity    

2. What is the before tax equilibrium price    

3. If the government imposes a per unit tax on each item, assuming the producers pay the tax, what is the new equilibrium quantity if the per unit tax is 3?    

4. Given the change in part 3, what is the new equilibrium price    


5. What is the price paid by the consumer   

6. What is the price received by the producer    

7. What is the per unit dollar value of the consumer's tax burden

    
8. What is the per unit dollar value of the producer's tax burden   
Answer #1

P = 24 -6Q & Demend Equation ③ P= 48 Suffly Equation =@ so Equating both 0 - 0 We get 24-60 = 40 .24-108 ..Q=24 1o 0=2:41 Solod=27 To-2.1 whits 4 So. P = 4(2.7)+3 | .:P 8.4+3 1P=$11.47 hit alt Rest del farts as a saferate questions -